OpenStudy (anonymous):

the sum of the units digit and 3 times the tens digit of a two-digit number is 14, and the sum of the tens digit and twice the units digit is 18. find the number.

OpenStudy (anonymous):

so Messi was in Bogota...

OpenStudy (anonymous):

@thewildones i know. how cool is that :)

OpenStudy (anonymous):

Let xy be the number. Then from the problem's restrictions:\[\{y+3x=14,x+2y=18\}\]Solve the above for x and y.\[\{x=2,y=8\} \]The number is:\[10*2+8=28 \]

OpenStudy (anonymous):

@robtobey thanks!

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

np!